OpenStudy (anonymous):

a coin purse contain a mixture of 32 coins in dimes and quarters. The coins have a total value of $5 or 500 cents. Determime. the number of dimes and the number of quarters in the purse.

OpenStudy (anonymous):

Let the number of dimes be x Let the number of quarters be y x+y=32 10x + 25y = 500 Solve for x and y x = 20 y= 12 Check, 20+12 = 32 20(10) + 12(25) = 500 Hence Proved.

OpenStudy (amistre64):

n + q = 32 n.05 + q.25 = 5 q = (5-32(.05))/(.25-.05) = 17 17(.25) + 15(.05) = 5.00

OpenStudy (amistre64):

why did i read nickels instead of dimes?

OpenStudy (anonymous):

Its 25c and 10c

OpenStudy (amistre64):

yeah, i see that now lol

OpenStudy (amistre64):

q = (5-32(.10))/(.25-.10) = 12 and the rest is yada yada ....

OpenStudy (anonymous):

thanks for both of your help.

OpenStudy (amistre64):

youre welcome, now if you ever need to know the one about the nickels ... lol

OpenStudy (anonymous):

I'm little lost with this equation q = (5-32(.10))/(.25-.10) = 12

OpenStudy (anonymous):

Keep it in cents, that way you don't have to deal with too many decimal points.

OpenStudy (amistre64):

its an application of cramers rule to determine the amount of quarters

OpenStudy (amistre64):

once we got the equations: d + q = 32 d.10 + q.25 = 5 we can eliminate the ds or qs; lets get rid of the ds by multiplying by -.10 d(-.10) + q(-.10) = 32(-.10) d.10 + q.25 = 5 ---------------------------- q(.25-.10) = 5-32(.10) ; and solve for q q= (5-32(.10))/(.25-.10)